The electrostatic force between two ions carrying equal charge is3.7 X 10-9 N .How many electrns are missing from each ion?

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
The electrostatic force b/w the two ions is given as:
F = kq^2/r^2

3.7 X 10 -9 N = 9 x 10 9q^2/r^2
0.4* 10 -18 = q
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)